OpenStudy (anonymous):

Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x2 + 13x + 15 2x2 + 27x + 13 5x2 − 2x − 7 10x2 − 7x − 3 8x2 − 11x + 3

OpenStudy (campbell_st):

pick the 1st one... so for a quadratic \[ax^2 + bx + c = 0\] multiply a and c so for \[2x^2 + 13x + 15 = 0\] 2 * 15 = 30 next find the factors of 30 that add to 30 10 + 3 so then write (ax + factor 1)(ax + factor 2) ----------------------------- =0 2 (2x + 10)(2x + 3) ---------------- =0 2 now look for common factors in either binomial so you get 2(x + 5)(2x + 3) ----------------- = 0 2 cancel the common factor of 2 leaves (x + 5)(2x + 3) = 0 now solve x + 5 = 0 and 2x + 3 = 0 for your solution.